AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 396     Entry time: Wed May 10 19:06:18 2023
Author: RSS, TD 
Subject: V-I test 

To test the detectors, adaptors were installed on the feed through flanges. A negative bias voltage was applied in steps of 10 V from 0 to 150 V. The V-I plots are attached for the four detectors. Three detectors - top right, top left, and bottom left (in the direction of the beam) - are fine. For the bottom right detector (in the direction of the beam), the current was fluctuating and was considerably high (1st test). The adaptors connecting to this detector were removed and checked by applying the bias voltage which showed zero current for voltages ranging from 0 to - 150 V, confirming the adaptors to be fine. They were then exchanged by other adaptors and the V-I test was repeated again (2nd test) with fluctuating and high current values. One of the flanges connecting to this detector was found to be light-tight. These tests were done at an ambient temperature of 21 degrees. Also, we couldn't find the rest of the three SHV to 2x Lemo 00 cables for the detector HV.

Attachment 1: detector_1.png  17 kB  Uploaded Wed May 10 20:06:44 2023  | Hide | Hide all
detector_1.png
Attachment 2: detector_2.png  18 kB  Uploaded Wed May 10 20:06:54 2023  | Hide | Hide all
detector_2.png
Attachment 3: detector_3.png  22 kB  Uploaded Wed May 10 20:07:01 2023  | Hide | Hide all
detector_3.png
Attachment 4: detector_4.png  18 kB  Uploaded Wed May 10 20:07:15 2023  | Hide | Hide all
detector_4.png
ELOG V3.1.4-unknown